i have to write an equation of a line that contains the given information:through the point (4,3) and parallel to the line y=2x+5
Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

3=2*4+b
3=8+b
b=-5
--------------
y=2x-5

Answer by stanbon(75887)

write an equation of a line that contains the given information:through the point (4,3) and parallel to the line y=2x+5
-------
If parallel to y = 2x+5, the slope is 2
---
Find equation equation of line with m = 2 and y = 3 when x = 4.
-------------------
Form: y = xm+b
Solve for "b:
3 = 2*4 + b
b = -5
-----
Equation:
y = 2x-5
==================
Cheers,
Stan H.
==================